OpenStudy (anonymous):

Find 3 consecutive intergers such that the product of the smallest and the largest is one less than nine times the middle interger.

OpenStudy (ujjwal):

let the numbers be x-1, x, x+1 so according to qsn (x-1)*(x+1)=9x-1 by solving it the numbers you get are 8,9, 10...

OpenStudy (anonymous):

let the 3 consecutive numbers equal x, x+1 and x+2, respectively. It follows (from the question) that (x)(x+2) = 9(x+1) - 1 therefore, \[x ^{2} -7x-8=0\] therefore, x = 8 or -1

OpenStudy (ujjwal):

yeah -1, 0, 1 also satisfies the solution.. :)